OpenStudy (anonymous):

FACTORING! please help me factor x^2+3x+9/4

OpenStudy (anonymous):

@jagr2713

myininaya (myininaya):

hint: this may help" \[x^2+bx+\frac{c}{a} \\ \frac{1}{a}(ax^2+bax+\frac{c(a)}{a}) \\ \frac{1}{a}(ax^2+bax+c)\]

OpenStudy (anonymous):

ok so my equation should look like \[1/4(4x^2+3(4)x+9)\]

myininaya (myininaya):

yep it is like dividing everything by 4 at the same time as multiplying everything by 4 and we know 4/4=1 so yep

myininaya (myininaya):

and 3(4) can be written as 12 of course

myininaya (myininaya):

you can also see if you need to factor anymore

OpenStudy (anonymous):

right, so now my equation should be \[x^2+12+9\]

myininaya (myininaya):

well you are missing 1/4( ) and a 4 in front of x^2 and a x next to the 12

myininaya (myininaya):

\[\frac{1}{4}(4x^2+12x+9)\]

myininaya (myininaya):

anyways let's look at the thing in ( )

myininaya (myininaya):

can you find two numbers that multiply to be 4(9)=36 and add up to be 12

OpenStudy (anonymous):

something that are multiples to 36 and add to 12?

myininaya (myininaya):

yep

OpenStudy (anonymous):

6 and 6

myininaya (myininaya):

yes yes! :)

myininaya (myininaya):

\[\frac{1}{4}(4x^2+6x+6x+9)\] now we can ignore the 1/4 and bring it down later but we need to factor the thing inside the ( )by grouping

myininaya (myininaya):

we will bring the 1/4 down later \[4x^2+6x+6x+9 \\ (4x^2+6x)+(6x+9) \\ 2x(2x+3)+(6x+9)\] you try factoring the 6x+9

OpenStudy (anonymous):

(6x+9) will equal 3x(2x+3)

myininaya (myininaya):

well 9 doesn't have a factor of x so you cannot factor an x out of it unless you divide by x but that is making is nastier than it should be 6x+9=3(2x+3)

myininaya (myininaya):

we will bring down 1/4 later \[2x(2x+3)+3(2x+3)\] now one last step

OpenStudy (anonymous):

oh ok, that makes sense!

OpenStudy (anonymous):

I'm not sure how to do the last step

myininaya (myininaya):

example to factor a(b+c)+m(b+c) notice both terms both a(b+c) and m(b+c) have the common factor (b+c) so I can factor (b+c) out of both terms in the sum (b+c)(a+m)

myininaya (myininaya):

so you see (2x+3) in both terms the terms being 2x(2x+3) and 3(2x+3)

myininaya (myininaya):

what is left of the first term when factoring out the (2x+3) is 2x and of the second term is +3

myininaya (myininaya):

2x(2x+3)+3(2x+3) (2x+3)(2x+3)

myininaya (myininaya):

don't forget to bring down the 1/4

OpenStudy (anonymous):

so its just 1/4(2x+3)(2x+3)

myininaya (myininaya):

or 1/4(2x+3)^2

myininaya (myininaya):

or...

myininaya (myininaya):

\[\frac{1}{4}(2x+3)^2 \\ \frac{1}{2^2}(2x+3)^2 \\ \frac{1^2}{2^2}(2x+3)^2 \\ (\frac{1}{2}(2x+3))^2 \\ (\frac{2x+3}{2})^2 \] there is another or :p

myininaya (myininaya):

you could separate the fraction inside the ( )^2

myininaya (myininaya):

but anyways yeah you are done

OpenStudy (anonymous):

so its just \[(x+\frac{ 3 }{ 2 })^2\]

myininaya (myininaya):

yep

OpenStudy (anonymous):

Thank you so much, I really appreciate your help!

myininaya (myininaya):

if we would have noticed the following to begin with: \[x^2+3x+(\frac{3}{2})^2=(x+\frac{3}{2})^2 \] we wouldn't had so much work

myininaya (myininaya):

\[x^2+kx+(\frac{k}{2})^2=(x+\frac{k}{2})^2 \]

OpenStudy (anonymous):

oh yes! That's very true, we missed that. Oh well :)

myininaya (myininaya):

for example: \[x^2+5x+(\frac{5}{2})^2=(x+\frac{5}{2})^2 \\ x^2-5x+(\frac{5}{2})^2=(x-\frac{5}{2})^2 \\ x^2-3x+(\frac{-3}{2})^2=(x-\frac{3}{2})^2 \\ x^2-2x+(\frac{-2}{2})^2=(x-1)^2 \]

OpenStudy (anonymous):

that actually makes perfect sense, thank you so much for showing me that formula! It's a lot easier than all the other ways I have been taught.

myininaya (myininaya):

no problem :)